OpenStudy (anonymous):

What is the sum of a 7-term geometric series if the first term is −6, the last term is −24,576, and the common ratio is −4?

OpenStudy (phantomcrow):

Do you know the formula?

OpenStudy (anonymous):

nope...

OpenStudy (anonymous):

@Deadly_Roses

OpenStudy (phantomcrow):

The formula for the sum of a geo series is:\[S _{n}=\frac{ a _{1} (1-r^n)}{ 1-r }\]

OpenStudy (anonymous):

would it be -34,766?

OpenStudy (phantomcrow):

That's not what I got. n=7 r=-4

OpenStudy (phantomcrow):

So\[S _{n}=-6(1-(-4)^7/1-(-4)\]

OpenStudy (anonymous):

is the fraction bar localized to 14^7/1, or does it encompass everything in parenthesis?

OpenStudy (phantomcrow):

everything

OpenStudy (anonymous):

im fried...

OpenStudy (phantomcrow):

Alright.\[S _{n}=\frac{ -6(1-(-4)^7) }{ 1-(-4)}\] This further simplifies to\[S_{n}=\frac{ -6(1-(-16834)) }{ 5 }\] Simplify again...\[S_{n}=\frac{ -6(1+16834) }{ 5 }\]\[S_{n}=\frac{ -98310 }{ 5 }\]\[S_{n}=-19662\]

OpenStudy (anonymous):

mind blown 0.0

OpenStudy (phantomcrow):

Haha. As long as you know the series formulas, you should be good to go for any type of these problems.

OpenStudy (anonymous):

alright. im gonna leave this open so i can write down the formula when i come back on Wednesday. thank you for your time and help

OpenStudy (phantomcrow):

No problem.